Question 11 (1 point)

The Prime Minister of Richlandia hires you as an economic consultant. He is concerned that the output level in Richlandia is too low and that this will cause unemployment to rise. He feels that it is necessary to increase output by $20 billion. He tells you that the MPC = 0.75. Which of the following would be the best advice to give to the Richlandia Prime Minister?

Question 11 options:

increase government purchases by $15 billion

reduce taxes by $10 billion

reduce taxes by $4 billion

increase government purchases by $5 billion

Question 12 (1 point)

Which of the following statements is correct?
I. The deficit and debt amounts must be presented as a percentage of GDP in order to have any meaning.
II. The government borrows money through issuing shares of stock.

Question 12 options:

I only

II only

both I and II

neither I nor II

Question 13 (1 point)

A minimal guideline for national debt is

Question 13 options:

the debt should grow no faster than nominal GDP

the debt should grow no faster than the budget deficit

the debt should grow no faster than the decrease in the tax rate

the debt should grow faster than nominal GDP
Answer #1

Question11

MPC = 0.75

Government Expenditure Multiplier = delta Y/deltaG = 1/1-MPC

Government Expenditure Multiplier = delta Y/deltaG = 1/1-0.75

Government Expenditure Multiplier = delta Y/deltaG = 4

delta Y/deltaG = 4

delta Y = 20

20/deltaG = 4

delta G = $5 billion

Also,

Tax Multiplier = deltaY/delta T = -MPC / 1-MPC

Tax Multiplier = deltaY/delta T = -0.75 / 1-0.75

Tax Multiplier = deltaY/delta T = - 3

20 / delta T = -3

delta T = -20/3 = -$6.67 billion

Since, delta G = $5billion

The government should increase government purchases by $5billion

Hence, 4th option is correct.

Question 12

Only statement I is correct

If the nominal values of deficit and debt are observed, then they can give a false impression about the country's GDP. They should be looked at as a percentage of GDP to see the relativity and actual situation of the economy. So, this statement is correct.

The government does not borrow money through the issue of stock. Stocks are issued by the stock market. The government generally borrows money by selling the government securities like bonds, treasury bills etc. So, this statement is wrong

Question 13

The first option is correct.

The national debt should grow no faster than nominal GDP.

If debt grows faster than nominal GDP then the overall debt to GDP ratio will increase. As a result, the interest payment on the national debt as a percentage of GDP will increase. The tax burden on the public will be higher.